A particle moves along x axis the velocity of particle as fiction of x is v²=12x-3x(in si).Find v at instant a=1.5

Grade:11

1 Answers

Vikas TU
6 years ago
Dear Student,
From the given data,V2 = 12x-3x = 9x
On differentiating both sides
=>2vdv=9dx
=>dv/dx=9/2v
=>a=9/2v (Since dv/dx=acceleration)
=>v=9/2a
=>v=3m/s
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)